# Optimising Communication Control Factors for Energy Consumption in Rural V2X

### v2

@@ II-A Traffic Model and Critical Distance
We consider a traffic consisting of a road side unit (RSU), an ego vehicle (Ego) and a target vehicle (Target) in a rural area. All the vehicles are registered with RSU, as illustrated in Fig. 1. Considering a scenario where Target is undergoing a hazard, as a result, RSU notified Ego by sending a brake signal via the communication channel. In this context, RSU transmits data packets containing brake signals, and Ego needs to receive at least one correct data packet to perform the brake. Error packets received by Ego causes a delay due to retransmission. Since the time RSU sent out the brake signals, Ego travels a distance until it first received a correct signal. This distance is known as a critical distance, e.g., $D_{\mathrm{comm}}$. To explain, $D_{\mathrm{comm}}$ demonstrates the distance that Ego travels during the time required by signal retransmission due to error, as illustrated in Fig. 2. At location $i=0$, RSU transmits a safety message in terms of packet to Ego periodically. Ego receives these packets sequentially at location indices $i=1,2,3,\ldots$. Due to channel errors, the process continues until the first error-free packet is successfully received at index $i=n$, which then provides the valid brake signal. During this interval, Ego continues to travel at its current speed, covering a distance referred to as $D_{\mathrm{comm}}$. In this context, $D_{\mathrm{comm}}$ can be considered as a safety threshold distance. The connected braking can execute successfully only when the distance between RSU and Ego is larger than $D_{\mathrm{comm}}$.
@@ II-B Simulation Setup
Scenario Rural highway, $2$ lanes, each $2000$ $\mathrm{m}$ long
@@
Mobility pattern Gaussian distributed speed, average $v$ and std. dev. $7$ km/h [WiLab2]
@@
Propagation model 3GPP Rural [3GPP]
Shadowing Variance 33 $\mathrm{dB}$, decorrelation distance $25$ $\mathrm{m}$ [WiLab1]
@@
Hybrid Automatic Repeat Request Blind retransmissions, max. $2$ allowed
@@
Subchannel size $10$ Physical Resource BlocK (PRB)
@@
Sensing threshold -$110$ $\mathrm{dBm}$
The V2X system in our rural scenario, considering traffic complexity and communication reliability, is simulated using WiLabV2XSim [WiLab1], an open source simulator that is well known for its accurate modelling of vehicular mobility, wireless channel dynamics, and communication reliability under varying traffic conditions. It provides a flexible and extensible framework for modelling V2X networks under various communication protocols, channel conditions, and mobility patterns [WiLab2, WiLab3].
@@
The channel model setting is generated following the 3GPP rural [3GPP]. Various $\mathrm{MCS}$ values are considered in the simulation: $\mathrm{MCS}$ 88 to $10$ represents the Quadrature Phase Shift Keying Modulation (QPSK), and $\mathrm{MCS}$ $12$ to $18$ perform the $16$-Quadrature Amplitude Modulation ($16$ QAM). In addition, $\mathrm{SCS}$ is modified between $15$ and $30$ $\mathrm{kHz}$ and ${P}_{\mathrm{t}}$ is modified from $23$ to $26$ $\mathrm{dBm}$. Different vehicle densities cause varying channel congestion, impacting delay and $\mathrm{PRR}$, and thereby altering the optimisation of communication control factors. In this work, we also consider shadowing. The value of shadowing $S_{i}$ is updated to its new value $S_{i+1}$ following the distance of the vehicle traveling as follows [3gpp36885]:
@@
where $d$ $\mathrm{m}\mathrm{e}\mathrm{t}\mathrm{e}\mathrm{r}\mathrm{s}$] is the change in distance between the RSU and the Ego. $d_{\mathrm{corr}}$ denotes the decorrelation distance of $25$ $\mathrm{m}$. Beyond $d_{\mathrm{corr}}$ the shadowing becomes increasingly uncorrelated and varies more independently. In addition, $N_{i+1}$ stands for a log normal independent random variable with standard deviation $\sigma=3$. We also use the free space pathloss (PL) model specified in 3GPP [3GPP]:
@@ III-B The Analysis of Critical Distance
Packets are transmitted periodically at rate $\mathrm{pps}$ packets per second, so each transmission slot lasts $1/\mathrm{pps}$ seconds. Assume that each packet reception is an independent and identically distributed (i.i.d) Bernoulli trial with the same success probability, the Ego vehicle is assumed to travel a set distance during the packet transmission attempt of the first packet, the expected number of failed receptions over the next N1N-1 attempts is $(N-1)(1-\mathrm{PRR})$. Each failure delays the first correctly received packet by exactly one slot; hence the expected additional waiting time (relative to the first attempt) is expressed as
@@ III-C The Analysis of Total Energy
Lemma 2 (Total energy spent on transmissions with truncated HARQ).
Let $\mathrm{pps}$ be the packet generation rate [pkt/s], $T$ be the horizon length [s]. Hence, there are $N_{\mathrm{pkt}}=\mathrm{pps}\,T$ packets generated. In addition, $L_{\mathrm{bits}}$ is the payload length [bits], $R$ denotes data rate [bit/s] and $H\in\mathbb{N}$ is the maximum number of attempts per generated packet (blind HARQ cap). Then, the expected total transmit energy over the horizon is
@@ IV Simulation Results
This section provides the simulation results, which are progressively structured to build on each other’s insights. We begin by analysing $D_{\mathrm{comm}}$ to determine the safety-critical communication threshold for the upcoming simulation traffic scenarios. Subsequently, the next simulation evaluates the effectiveness of different $\mathrm{SCS}$ settings. Based on this, an analysis of $P_{t}$ under different traffic conditions is investigated. The next simulation focuses on analysing $\mathrm{MCS}$ effectiveness while the final simulation campaign explores the energy consumption analysis, incorporating optimised communication control factors derived from the findings of the previous simulations.
@@ IV-D The Impact of MCS on PRR
TABLE III: Maximum total energy cost ($J$) for delivering packets from RSU to all vehicles at $\mathrm{SCS}=30$, $\mathrm{MCS}=8$, and $v$ = $50$ $\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$.
@@ V Analysis on Energy Consumption
This section provides the analysis of energy consumption, which enables us to optimise the $\mathrm{MCS}$ configuration and the most suitable code rate. The target is to achieve a good balance between $\mathrm{PRR}$ performance and energy efficiency under both light and heavy traffic conditions.
@@ V-A The Impact of MCS on Achieved $E_{b}/N_{0}$
Fig. 13 and Fig. 14 show the obtained $E_{b}/N_{0}$ under various $\mathrm{MCS}$ configurations the light and heavy traffic, respectively. The effect of ${P}_{\mathrm{t}}$, ranging from $23$ to $26$ $\mathrm{dBm}$, is examined for vehicle densities of $50$ and $100$ vehicles per $\mathrm{km}$, revealing trade-offs between energy efficiency and communication reliability. As can be seen from Fig. 13, under light traffic of $50$ vehicles per $\mathrm{km}$, increasing ${P}_{\mathrm{t}}$ is expected to make significant improvements across all $\mathrm{MCS}$ values. However, noticeable gains are only observed at extremely high $\mathrm{MCS}$ levels, $16$ to $18$, which is not energy efficient. For lower $\mathrm{MCS}$ values of $8$ to $10$, the system maintains consistently robust performance, with $E_{b}/N_{0}$ values remaining above $10$ $\mathrm{dB}$ throughout the range of $D_{\mathrm{comm}}$, indicating reliable communication quality under the tested conditions. As shown in Fig. 13(a), within the $D_{\text{comm}}$, $\mathrm{MCS}$ 88 achieves an $E_{b}/N_{0}$ of $10.2$ $\mathrm{dB}$ at $23$ $\mathrm{dBm}$ and $10.5$ $\mathrm{dB}$ at $26$ $\mathrm{dBm}$, indicating only a marginal improvement. Compare with Fig. 13(b), these findings suggest that the use of $\mathrm{MCS}$ 88 to $10$ with the setting ${P}_{\mathrm{t}}=23$ $\mathrm{dBm}$ is energy efficient and sufficient for light traffic scenarios. In contrast, the worst case is shown in Fig. 14 which exhibits a tighter clustering of all $\mathrm{MCS}$ curves and a reduced average $E_{b}/N_{0}$. This indicates higher interference and reduced link quality. Moreover, compare between the results shown in Fig. 14(a) and 14(b), the performance gap between $26$ and $23$ $\mathrm{dBm}$ reduces, and the benefit of using higher $\mathrm{MCS}$ becomes less pronounced. Under such conditions, lower $\mathrm{MCS}$ levels, such as $8$ to $10$ become more resilient and energy efficient. Higher $\mathrm{MCS}$ levels are struggling to maintain sufficient $E_{b}/N_{0}$ beyond $300$ m, which results in increased retransmissions and higher energy consumption. Finally, based on the analysis under both light and heavy traffic conditions, a lower level $\mathrm{MCS}$ with a value of $8$, which demonstrates consistently good performance across scenarios. Therefore, $\mathrm{MCS}$ 88 is selected to analyse the trade-offs between energy consumption and safety in the next subsections.
@@ V-B The Impact of $P_{t}$ on Achieved $E_{b}/N_{0}$
Fig. 15 shows the achieved $E_{b}/N_{0}$ against the distance between RSU and Ego, under various ${P}_{\mathrm{t}}$ values of $23$, $24$, $25$, and $26$ $\mathrm{dBm}$ and vehicle densities of $30$, $50$, $80$, and $100$ vehicles per $\mathrm{km}$. The corresponding energy expenditure, measured as the average number of transmissions required to successfully deliver $350$ packets, is summarised in Table II. These results provide insights on the configuration of ${P}_{\mathrm{t}}$ and $\mathrm{MCS}$ to minimise energy consumption while maintaining reliable V2X connectivity in ruralenvironments. According to Fig. 15, the results reveal a clear trade-off between energy efficiency and communication reliability across vehicle density scenarios. At low densities (e.g., $30$ to $50$ vehicles per $\mathrm{km}$), reducing ${P}_{\mathrm{t}}$ from $26$ to $23$ $\mathrm{dBm}$ leads to only minor reductions in $E_{b}/N_{0}$, particularly in the short to medium ranges. For example, the loss is approximately $1.5$ $\mathrm{dB}$ at $100$ $\mathrm{m}$ and remains within a tolerable margin at $200$ $\mathrm{m}$ for stable packet reception. These findings indicate that employing a lower ${P}_{\mathrm{t}}$, such as $23$ $\mathrm{dBm}$, is sufficient in low-density traffic to conserve energy without compromising communication reliability, thus improving the efficiency of V2X systems in rural scenarios.
@@ V-C Analysis on Energy Consumption
Table II illustrates the averaged total generated transmissions from RSU to all the vehicles, based on optimising ${P}_{\mathrm{t}}$. As shown in Table II, the total number of transmissions does not strictly decrease monotonically with $P_{\mathrm{t}}$. This behaviour is due to randomness in the simulation, caused by shadowing fading, traffic mobility, and random scheduling of retransmissions, all of which introduce small fluctuations in aggregate transmission counts. However, a general decreasing trend can still be observed in the results. Under light traffic conditions ($30$ and $50$ vehicles per $\mathrm{km}$), increasing $P_{\mathrm{t}}$ from $23$ to $25$ $\mathrm{d}\mathrm{B}\mathrm{m}$ decreased the total number of transmissions for about $30000$. A similar pattern is observed under heavy traffic ($30$ and $50$ vehicles per $\mathrm{km}$), where increasing $P_{\mathrm{t}}$ from $25$ to $26~\mathrm{dBm}$ provides moderate reliability improvements while reducing the retransmission overhead.
@@
Table III illustrates the maximum total energy for delivering all packets from RSU to all vehicles. Under light and moderate traffic densities of $30$ and $50$ vehicles per $\mathrm{km}$, the increase in ${P}_{\mathrm{t}}$ from $23$ to $26$ $\mathrm{dBm}$ results in a more gradual increase in total energy from $90.6$ to $157.52~\mathrm{J}$ and from $300.68$ to $579.09~\mathrm{J}$, representing approximately a 7474% and 9393% rise, respectively. This shows that in sparse environments, higher ${P}_{\mathrm{t}}$ can improve coverage and $\mathrm{PRR}$ with only a moderate increase in energy consumption. However, once traffic density exceeds a critical threshold (e.g. $80$ vehicles per $\mathrm{km}$), the marginal energy gain per successfully delivered packet decreases sharply. In contrast to light traffic conditions, where moderate increases in ${P}_{\mathrm{t}}$ yield energy-efficient gains, heavy traffic scenarios exhibit a different trade-off. At $\rho=100$, raising ${P}_{\mathrm{t}}$ from $23$ to $26$ $\mathrm{dBm}$ increases the total number of transmissions and consequently leads to a substantial increase in total energy consumption, from about $1255.92~\mathrm{J}$ to $2523.51~\mathrm{J}$, the growth in total energy cost is much more pronounced than light traffics, approximately 101101%. This steep growth indicates that in dense vehicular networks, additional ${P}_{\mathrm{t}}$ has proportionally improved communication reliability, but also amplified the overall energy cost to deliver all packets by RSU.
@@
The results highlight the trade-off between communication reliability and energy efficiency in V2X networks. Increasing ${P}_{\mathrm{t}}$ effectively extends $D_{\mathrm{comm}}$ and enhances the PRR, thus improving transmission reliability and road safety. However, this improvement shows inconsistency, which can be attributed to random fluctuations of shadowing and traffic layouts during simulations. Consequently, total energy consumption increases sharply with ${P}_{\mathrm{t}}$, in light traffic scenarios, a lower ${P}_{\mathrm{t}}$ of $23$ $\mathrm{dBm}$ is sufficient to maintain effective communication while reducing energy costs. However, in high-density scenarios, a higher ${P}_{\mathrm{t}}$ of $26$ $\mathrm{dBm}$ is effective in extending the scenario $D_{\mathrm{comm}}$ to improve safety performance, but this cost of energy from excessive power needs to be considered, as it does not effectively reduce the total number of transmissions. These findings suggest that, while increasing ${P}_{\mathrm{t}}$ improves connectivity and safety performance, it does so at the expense of energy efficiency. Hence, a balance must be achieved between reliable communication and sustainable energy use. Therefore, an adaptive power control mechanism that dynamically adjusts ${P}_{\mathrm{t}}$ according to vehicle density is essential to optimise both safety performance and overall energy consumption in future rural V2X systems.

### v3

@@ II-A Traffic Model and Critical Distance
We consider a traffic consisting of a road side unit (RSU) [^1] In our rural scenario, RSU is assumed to operate as a communication-centric V2X infrastructure node, focusing on sidelink communication support rather than environmental perception or sensor-based detection., an ego vehicle (Ego) and a target vehicle (Target) in a rural area. All the vehicles are registered with RSU, as illustrated in Fig. 1. Considering a scenario where Target is undergoing a hazard, as a result, RSU notified Ego by sending a brake signal via the communication channel. In this context, RSU transmits data packets containing brake signals, and Ego needs to receive at least one correct data packet to perform the brake. Error packets received by Ego causes a delay due to retransmission. Since the time RSU sent out the brake signals, Ego travels a distance until it first received a correct signal. This distance is known as a critical distance, e.g., $D_{\mathrm{comm}}$. To explain, $D_{\mathrm{comm}}$ demonstrates the distance that Ego travels during the time required by signal retransmission due to error, as illustrated in Fig. 2. At location $i=0$, RSU transmits a safety message in terms of packet to Ego periodically. Ego receives these packets sequentially at location indices $i=1,2,3,\ldots$. Due to channel errors, the process continues until the first error-free packet is successfully received at index $i=n$, which then provides the valid brake signal. During this interval, Ego continues to travel at its current speed, covering a distance referred to as $D_{\mathrm{comm}}$. In this context, $D_{\mathrm{comm}}$ can be considered as a safety threshold distance. The connected braking can execute successfully only when the distance between RSU and Ego is larger than $D_{\mathrm{comm}}$.
@@ II-B Simulation Setup
Scenario 22 lanes rural highway, $2000$ $\mathrm{m}$
@@
Mobility pattern ($V$) $V\sim\mathcal{N}(v,\,7^{2})\ \text{km/h}$ [5]
@@
Propagation model 3GPP Rural [3]
Shadowing Variance 33 $\mathrm{dB}$, decorrelation distance $25$ $\mathrm{m}$ [23]
@@
Hybrid Automatic Repeat Request Blind retransmissions, max. 22
Average circuit power per transmission $(P_{c})$ $100$ $\mathrm{mW}$ [34]
Average idle power $(P_{i})$ $125$ $\mathrm{mW}$ [16]
@@
Subchannel size $10$ Physical Resource Blocks (PRB)
@@
Sensing threshold 110–110 $\mathrm{dBm}$
The V2X system in our rural scenario, considering traffic complexity and communication reliability, is simulated using WiLabV2XSim [23], an open source simulator that is well known for its accurate modelling of vehicular mobility, wireless channel dynamics, and communication reliability under varying traffic conditions. It provides a flexible and extensible framework for modelling V2X networks under various communication protocols, channel conditions, and mobility patterns [5, 11].
@@
The channel model setting is generated following the 3GPP rural [3]. Various $\mathrm{MCS}$ values are considered in the simulation: $\mathrm{MCS}$ 88 to $10$ represents the Quadrature Phase Shift Keying Modulation (QPSK), and $\mathrm{MCS}$ $12$ to $18$ perform the $16$-Quadrature Amplitude Modulation ($16$ QAM). In addition, $\mathrm{SCS}$ is modified between $15$ and $30$ $\mathrm{kHz}$ and ${P}_{\mathrm{t}}$ is modified from $23$ to $26$ $\mathrm{dBm}$. During the sensing window, RSU measures the received signal reference power (RSRP) associated with decoded sidelink control information (SCI), and the sensing threshold is set to 110-110 $\mathrm{d}\mathrm{B}\mathrm{m}$ in this study. Different vehicle densities cause varying channel congestion, impacting delay and $\mathrm{PRR}$, and thereby altering the performance evaluation of communication control factors. In this work, we also consider shadowing. The value of shadowing $S_{i}$ is updated to its new value $S_{i+1}$ following the distance of the vehicle traveling as follows [1]:
@@
where $d$ $\mathrm{m}\mathrm{e}\mathrm{t}\mathrm{e}\mathrm{r}\mathrm{s}$] is the change in distance between the RSU and the Ego. $d_{\mathrm{corr}}$ denotes the decorrelation distance of $25$ $\mathrm{m}$. Beyond $d_{\mathrm{corr}}$ the shadowing becomes increasingly uncorrelated and varies more independently. In addition, $N_{i+1}$ stands for a log normal independent random variable with standard deviation $\sigma=3$. We also use the free space pathloss (PL) model specified in 3GPP [3]:
@@ III-B The Analysis of Critical Distance
Packets are transmitted periodically at rate $\mathrm{pps}$ packets per second, so each transmission slot lasts $1/\mathrm{pps}$ seconds. In this work, for tractable analysis, we assume that each packet reception is an independent and identically distributed (i.i.d) Bernoulli trial with the same success probability [2], the Ego vehicle is assumed to travel a set distance during the packet transmission attempt of the first packet, the expected number of failed receptions over the next N1N-1 attempts is $(N-1)(1-\mathrm{PRR})$. Each failure delays the first correctly received packet by exactly one slot; hence the expected additional waiting time (relative to the first attempt) is
@@ III-C The Analysis of Total Energy
Lemma 2 (Total energy spent on RF transmissions with truncated HARQ).
Let $\mathrm{pps}$ be the packet generation rate [pkt/s], $T$ be the horizon length [s]. Hence, there are $N_{\mathrm{pkt}}=\mathrm{pps}\,T$ packets generated. In addition, $L_{\mathrm{bits}}$ is the payload length [bits], $R$ denotes data rate [bit/s] and $H\in\mathbb{N}$ is the maximum number of attempts per generated packet (blind HARQ cap). Then, the expected total RF transmit energy over the horizon is
@@
Lemma 3 (Total communication energy).
let $P_{\mathrm{c}}$ denote the additional averaged circuit power consumed during transmission. Let $P_{\mathrm{i}}$ denote the averaged platform idle power over the horizon $T$. Then the expected total communication energy over the horizon is
$\!\!\!E_{\mathrm{comm}}$ $=N_{\mathrm{pkt}}\!\left[(P_{t}+P_{c})\frac{L_{\mathrm{bits}}}{R}\frac{1-(1-\mathrm{PRR})^{H}}{\mathrm{PRR}}+\frac{P_{\mathrm{i}}}{\mathrm{pps}}\right].$
Proof.
Based on Lemma 2, during each attempt, an averaged circuit power $P_{c}$ is associated with the active transmission. Therefore, the expected communication energy spent on transmissions for one generated packet is
$E_{\mathrm{pkt,tx}}=(P_{t}+P_{c})\,\frac{L_{\mathrm{bits}}}{R}\,\frac{1-(1-\mathrm{PRR})^{H}}{\mathrm{PRR}},$ (14)
Next, the corresponding idle energy over the horizon is $E_{\mathrm{idle}}=P_{i}\,T$, and in per-packet allocation:
$E_{\mathrm{idle}}=P_{i}\,T=N_{\mathrm{pkt}}\;\frac{P_{i}}{\mathrm{pps}},$ (15)
Combing  (14) and  (15) gives us
$E_{\mathrm{comm}}$ $=N_{\mathrm{pkt}}\left(E_{\mathrm{pkt,tx}}+\frac{P_{i}}{\mathrm{pps}}\right).$ (16)
which claims the lemma. ∎
@@ IV Simulation Results
This section provides the simulation results, which are progressively structured to build on each other’s insights. We begin by analysing $D_{\mathrm{comm}}$ to determine the safety-critical communication threshold for the upcoming simulation traffic scenarios. Subsequently, the next simulation evaluates the effectiveness of different $\mathrm{SCS}$ settings. Based on this, an analysis of $P_{t}$ under different traffic conditions is investigated. The next simulation focuses on analysing $\mathrm{MCS}$ effectiveness while the final simulation campaign explores the RF energy consumption analysis, incorporating the performance evaluation of communication control factors derived from the findings of the previous simulations.
@@ IV-D The Impact of MCS on PRR
TABLE III: Maximum RF energy cost ($J$) for delivering packets from RSU to all vehicles at $\mathrm{SCS}=30$, $\mathrm{MCS}=8$, and $v$ = $50$ $\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$.
@@
TABLE IV: Maximum communication energy cost ($J$) for delivering packets from RSU to all vehicles at $\mathrm{SCS}=30$, $\mathrm{MCS}=8$, and $v$ = $50$ $\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$.
$P_{\mathrm{t}}$ ($\mathrm{dBm}$) $\rho=30$ $\rho=50$ $\rho=80$ $\rho=100$
$23$ $140.52$ $455.75$ $1231.70$ $1889.75$
$24$ $148.30$ $498.68$ $1407.17$ $2244.56$
$25$ $164.05$ $596.95$ $1687.16$ $2675.75$
$26$ $201.46$ $728.93$ $2250.75$ $3161.76$
@@ V Analysis on Energy Consumption
This section provides the analysis of RF energy consumption, which enables us to evaluate the impact of $\mathrm{MCS}$ configuration and identify the most suitable code rate. The target is to achieve a good balance between $\mathrm{PRR}$ performance and energy efficiency under both light and heavy traffic conditions.
@@ V-A The Impact of MCS on Achieved $E_{b}/N_{0}$
Fig. 13 and Fig. 14 show the obtained $E_{b}/N_{0}$ under various $\mathrm{MCS}$ configurations the light and heavy traffic, respectively. The effect of ${P}_{\mathrm{t}}$, ranging from $23$ to $26$ $\mathrm{dBm}$, is examined for vehicle densities of $50$ and $100$ vehicles per $\mathrm{km}$, revealing trade-offs between energy efficiency and communication reliability. As can be seen from Fig. 13, under light traffic of $50$ vehicles per $\mathrm{km}$, increasing ${P}_{\mathrm{t}}$ is expected to make significant improvements across all $\mathrm{MCS}$ values. However, noticeable gains are only observed at extremely high $\mathrm{MCS}$ levels, $16$ to $18$, which is not energy efficient. For lower $\mathrm{MCS}$ values of $8$ to $10$, the system maintains consistently robust performance, with $E_{b}/N_{0}$ values remaining above $10$ $\mathrm{dB}$ throughout the range of $D_{\mathrm{comm}}$, indicating reliable communication quality under the tested conditions. As shown in Fig. 13(a), within the $D_{\text{comm}}$, $\mathrm{MCS}$ 88 achieves an $E_{b}/N_{0}$ of $10.2$ $\mathrm{dB}$ at $23$ $\mathrm{dBm}$ and $10.5$ $\mathrm{dB}$ at $26$ $\mathrm{dBm}$, indicating only a marginal improvement. Compare with Fig. 13(b), these findings suggest that the use of $\mathrm{MCS}$ 88 to $10$ with the setting ${P}_{\mathrm{t}}=23$ $\mathrm{dBm}$ is energy efficient and sufficient for light traffic scenarios. In contrast, the worst case is shown in Fig. 14 which exhibits a tighter clustering of all $\mathrm{MCS}$ curves and a reduced average $E_{b}/N_{0}$. This indicates higher interference and reduced link quality. Moreover, compare between the results shown in Fig. 14(a) and 14(b), the performance gap between $26$ and $23$ $\mathrm{dBm}$ reduces, and the benefit of using higher $\mathrm{MCS}$ becomes less pronounced. Under such conditions, lower $\mathrm{MCS}$ levels, such as $8$ to $10$ become more resilient and energy efficient. Higher $\mathrm{MCS}$ levels are struggling to maintain sufficient $E_{b}/N_{0}$ beyond $300$ m, which results in increased retransmissions and higher RF energy consumption. Finally, based on the analysis under both light and heavy traffic conditions, a lower level $\mathrm{MCS}$ with a value of $8$, which demonstrates consistently good performance across scenarios. Therefore, $\mathrm{MCS}$ 88 is selected to analyse the trade-offs between RF energy consumption and safety in the next subsections.
@@ V-B The Impact of $P_{t}$ on Achieved $E_{b}/N_{0}$
Fig. 15 shows the achieved $E_{b}/N_{0}$ against the distance between RSU and Ego, under various ${P}_{\mathrm{t}}$ values of $23$, $24$, $25$, and $26$ $\mathrm{dBm}$ and vehicle densities of $30$, $50$, $80$, and $100$ vehicles per $\mathrm{km}$. The corresponding energy expenditure, measured as the average number of transmissions required to successfully deliver $350$ packets, is summarised in Table II. These results provide insights on the configuration of ${P}_{\mathrm{t}}$ and $\mathrm{MCS}$ to minimise RF energy consumption while maintaining reliable V2X connectivity in ruralenvironments. According to Fig. 15, the results reveal a clear trade-off between energy efficiency and communication reliability across vehicle density scenarios. At low densities (e.g., $30$ to $50$ vehicles per $\mathrm{km}$), reducing ${P}_{\mathrm{t}}$ from $26$ to $23$ $\mathrm{dBm}$ leads to only minor reductions in $E_{b}/N_{0}$, particularly in the short to medium ranges. For example, the loss is approximately $1.5$ $\mathrm{dB}$ at $100$ $\mathrm{m}$ and remains within a tolerable margin at $200$ $\mathrm{m}$ for stable packet reception. These findings indicate that employing a lower ${P}_{\mathrm{t}}$, such as $23$ $\mathrm{dBm}$, is sufficient in low-density traffic to conserve energy without compromising communication reliability, thus improving the efficiency of V2X systems in rural scenarios.
@@ V-C Analysis on Energy Consumption
Table II illustrates the averaged total generated transmissions from RSU to all the vehicles, based on increasing ${P}_{\mathrm{t}}$. As shown in Table II, the total number of transmissions does not strictly decrease monotonically with $P_{\mathrm{t}}$. This behaviour is due to randomness in the simulation, caused by shadowing fading, traffic mobility, and random scheduling of retransmissions, all of which introduce small fluctuations in aggregate transmission counts. However, a general decreasing trend can still be observed in the results. Under light traffic conditions ($30$ and $50$ vehicles per $\mathrm{km}$), increasing $P_{\mathrm{t}}$ from $23$ to $25$ $\mathrm{d}\mathrm{B}\mathrm{m}$ decreased the total number of transmissions for about $30000$. A similar pattern is observed under heavy traffic ($30$ and $50$ vehicles per $\mathrm{km}$), where increasing $P_{\mathrm{t}}$ from $25$ to $26~\mathrm{dBm}$ provides moderate reliability improvements while reducing the retransmission overhead.
@@
Table III illustrates the maximum RF energy for delivering all packets from RSU to all vehicles. Under light and moderate traffic densities of $30$ and $50$ vehicles per $\mathrm{km}$, the increase in ${P}_{\mathrm{t}}$ from $23$ to $26$ $\mathrm{dBm}$ results in a more gradual increase in total energy from $90.6$ to $157.52~\mathrm{J}$ and from $300.68$ to $579.09~\mathrm{J}$, representing approximately a 7474% and 9393% rise, respectively. This shows that in sparse environments, higher ${P}_{\mathrm{t}}$ can improve coverage and $\mathrm{PRR}$ with only a moderate increase in RF energy consumption. However, once traffic density exceeds a critical threshold (e.g. $80$ vehicles per $\mathrm{km}$), the marginal energy gain per successfully delivered packet decreases sharply. In contrast to light traffic conditions, where moderate increases in ${P}_{\mathrm{t}}$ yield energy-efficient gains, heavy traffic scenarios exhibit a different trade-off. At $\rho=100$, raising ${P}_{\mathrm{t}}$ from $23$ to $26$ $\mathrm{dBm}$ increases the total number of transmissions and consequently leads to a substantial increase in total RF energy consumption, from about $1255.92~\mathrm{J}$ to $2523.51~\mathrm{J}$, the growth in total energy cost is much more pronounced than light traffics, approximately 101101%. This steep growth indicates that in dense vehicular networks, additional ${P}_{\mathrm{t}}$ has proportionally improved communication reliability, but also amplified the overall energy cost to deliver all packets by RSU. Table IV shows the total communication energy expense with various values of ${P}_{\mathrm{t}}$. The results show that RF energy still contributes more than $60\%$ of the overall communication energy across the considered configurations. This indicates that RF energy consumption remains a dominant and controllable component, and therefore motivates focusing on selecting appropriate communication control factors to reduce marginal RF energy while maintaining reliability in rural V2X scenarios.
@@
The results highlight the trade-off between communication reliability and energy efficiency in V2X networks. Increasing ${P}_{\mathrm{t}}$ effectively extends $D_{\mathrm{comm}}$ and enhances the PRR, thus improving transmission reliability and road safety. However, this improvement shows inconsistency, which can be attributed to random fluctuations of shadowing and traffic layouts during simulations. Consequently, total RF energy consumption increases sharply with ${P}_{\mathrm{t}}$, in light traffic scenarios, a lower ${P}_{\mathrm{t}}$ of $23$ $\mathrm{dBm}$ is sufficient to maintain effective communication while reducing energy costs. However, in high-density scenarios, a higher ${P}_{\mathrm{t}}$ of $26$ $\mathrm{dBm}$ is effective in extending the scenario $D_{\mathrm{comm}}$ to improve safety performance, but this cost of energy from excessive power needs to be considered, as it does not effectively reduce the total number of transmissions. These findings suggest that, while increasing ${P}_{\mathrm{t}}$ improves connectivity and safety performance, it does so at the expense of energy efficiency. Hence, a balance must be achieved between reliable communication and sustainable energy use. Therefore, an adaptive power control mechanism that dynamically adjusts ${P}_{\mathrm{t}}$ according to vehicle density is essential to improve both safety performance and overall RF energy efficiency in future rural V2X systems.